OpenStudy (anonymous):

find the value of x when 2(x-3) + 1 = 3x + 4

hartnn (hartnn):

are you stuck? at which step?

OpenStudy (anonymous):

2(x-3)+1=3x+4 2x-6+1=3x-4 2x-5-3x+4=0 -x-1 x=-1

OpenStudy (anonymous):

I had -1 and I got told it was wrong /:

OpenStudy (anonymous):

@Jbear* its like that -x =1

OpenStudy (anonymous):

My only answer choices are, -1 , -1 over 5, -9over 5, and -9.

OpenStudy (tamtoan):

it's wrong because of the mistake in during second step...3x +4 and rewrite 3x - 4... if you want to check answer right or wrong, just sub the value of x in and if you don't get both side equal, then it's not right.

hartnn (hartnn):

@Ekaansh please do not provide entire solution , involve the asker

hartnn (hartnn):

2(x-3)+1=3x+4 2x-6+1 = 3x+4 did u get this step ? @Jbear*

OpenStudy (anonymous):

Yes

hartnn (hartnn):

so what is -6+1 = ?

hartnn (hartnn):

2x-5=3x+4 Add 5 to both sides, what u get ?

hartnn (hartnn):

@Jbear* did u try to add 5 ?

hartnn (hartnn):

are you having any difficulty? feel free to express.....

hartnn (hartnn):

when u add 5 you get. 2x=3x+4+5 =3x+9 now subtract 3x from both sides, u get 2x-3x=9 -x=9 or x=-9 hope that you found your mistake. thanks.

OpenStudy (anonymous):

-9

OpenStudy (anonymous):

Answer is - 9